Multiply.
4xy^3*2x^2*5xy^2

Answer by AnlytcPhil(1116)

Multiply.

4xy3·2x2·5xy2

Put a 1 exponent on the letters that don't
show an exponent

4x1y3·2x2·5x1y2

Multiply the coefficients:  4·2·5 = 40

So far you have 40_______

Add the exponents of the x's:  1+2+1 = 4  

So far you have 40x4_____

Add the exponents of the y's: 3+2 = 5

Now you have 40x4y5

That's the answer.

Edwin McCravy
AnlytcPhil@aol.com